# Layered, Overlapping, and Inconsistent: A Large-Scale Analysis of the Multiple Privacy Policies and Controls of U.S. Banks

### v1

@@ 4.1. Data Collection
We used a list of the largest commercial banks by the Federal Reserve (of Governors of the Federal Reserve System, 2024), with consolidated assets exceeding \$300 million, each uniquely identified by its RSSD ID, a Federal Reserve-assigned identifier ($n$=2129). [^1] Consumer-oriented institutions also include savings banks and state non-member banks. They operate at smaller, regional scales and were not included in our study. We use the consolidated asset total as a proxy for customer base and consumer reach: these banks account for 97.3% of all assets held by FDIC-insured commercial banks (52.8% of them by number). We used the FDIC’s BankFind Suite (Federal Deposit Insurance Corporation, 2025) to obtain bank website URLs and branch locations based on their RSSD IDs. We removed 19 duplicate URLs associated with banks owned by the same holding company, [^2] A bank holding company may include multiple legally distinct banks (each with a unique RSSD ID) but typically provides a shared privacy policy across subsidiaries. as well as 19 duplicates where different URLs redirected to the same website due to mergers or acquisitions. We retained only the highest-ranked occurrence in each duplicate case, resulting in 2,091 unique bank website URLs. Because our study includes analysis of banks’ CCPA-related disclosures and opt-out implementations, we collected all data from a California vantage point using a commercial VPN. We successfully accessed 2,073 bank websites, which comprise the final dataset for this study. For each bank, we collected its (1) privacy policies, (2) cookie opt-out controls, (3), CCPA opt-out links, (4) responses to the Global Privacy Control (GPC) signal, and (5) examined third-party cookies. Data collection ran from October 2024 to January 2025. Figure  2 shows our data collection pipeline.
@@ 4.1.1. Privacy Policy Collection
Privacy Policy Retrieval
We built a custom crawler based on Scrapy (Developers, 2008) that searches for links containing a list of privacy policy-related keywords in either the link’s anchor text or URL. Our keyword list was informed by prior research (Amos et al., 2021; Libert, 2018; Wagner, 2023b; Srinath et al., 2021) and iteratively refined and tested on a sample of 100 banks. Our crawler searched for relevant links up to a depth of three to ensure reliable discovery of privacy policies. It successfully accessed 91.31% of webpages and successfully downloaded 100% of identified potential privacy policies. We manually visited 373 bank websites where the crawler failed to access pages, found no privacy-related content, or no GLBA notices (which we expected all banks to provide). We also manually reviewed the top 200 banks due to more complex site structures. In total, we collected 11,265 potential privacy policy files. Through automated and manual review, we filtered out non-relevant files and duplicates, resulting in 3,372 unique privacy files across 2,069 banks.
@@
Often banks had a webpage titled “Privacy Policy” that contained a collection of privacy policy information, such as a GLBA notice, a CCPA privacy policy and/or a CCPA notice at collection, data practices related to their online and mobile services, and/or a cookie policy. Some banks presented some or all notices in separate, stand-alone policies. To answer RQ1 (number of privacy policies a bank provides), we manually labeled files based on headings and categorized them into five types: (1) GLBA notice, typically titled “U.S. consumer privacy notice,” or “privacy policy.” It is typically presented as a PDF or sometimes as HTML, either standalone or combined with other notices and resembles the GLBA short-form template. (2) General privacy policy, often titled “privacy policy” or “digital privacy,” focused on online privacy, or may combine elements that would otherwise appear in the following three types of standalone notices: (3) CCPA privacy policy, by which we refer to both the privacy policy and the notice at collection provided in compliance with the CCPA; (4) mobile privacy policy, specific to mobile applications or mobile data collection practices; (5) cookie/advertising policy, focusing on cookie use and tracking technologies or interest-based advertising.
@@ 4.1.3. Third-Party cookies
We used third-party cookies as one indicator of whether a bank shares consumer information with external entities. We rendered each site using a Selenium-based web browser. We then used the Chrome DevTools Protocol’s Network domain to monitor network activity and collect all cookies stored during the browser session. Each cookie’s domain was compared to the domain of the website visited; if the domains did not match, the cookie was classified as a third-party cookie. After collecting third-party cookies from each website, we attempted to infer their purpose by comparing their domains against EasyList’s known list of advertising-related trackers (Fanboy et al., 2025). If no exact match was found, we compared only the last two segments of the cookie’s domain with the list (e.g., ”unrulymedia.com” instead of ”.targeting.unrulymedia.com”), which improved detecting advertising-related cookies. We found 1454 banks (70.1%) allowing third-party cookies on their websites, 1252 (60.4%) of which contained marketing cookies.
@@ 4.2. Privacy Policy Analysis
To address RQ1 on the amount of privacy information provided by each bank, we extracted the plain text for all of a bank’s policy files. For PDF files, we used a vision-based LLM pipeline (GPT-4o) to extract and compile text from each page (see Appendix LABEL:appendix:plain-text-extraction for details), as text extraction methods (e.g, using PyMuPDF (Inc., 2025)) proved unreliable due to the visual complexity of the layouts (e.g., flipped cell texts in GLBA notice tables). For HTML files, we used Boilerpipe (Riebold, 2023) to extract the main text and BeautifulSoup (Richardson, 2025) as a fallback. Then, we merged all of a bank’s processed privacy policy files into one plain-text file. For readability analysis, we used the widely-adopted Flesh-Kincaid Grade Level (Shedlosky-Shoemaker et al., 2009) that measures the difficulty of reading a text based on sentence length and word complexity with a score corresponding to a U.S. school grade (see Appendix LABEL:appendix:readability for results with other readability metrics).
@@ 4.2.2. GLBA Notice Analysis
GLBA notices mostly follow the standardized template, with the majority in PDF or HTML format. We analyzed the data-sharing table, the opt-out box, and the OII box in them (see Figure 1). For PDFs, we employed our vision-based LLM pipeline and refined it given the known GLBA table format: we first manually annotated the page numbers of each target item and then extracted the relevant items from each page into a prescribed format using task-specific prompts (see Figure LABEL:fig:glba-analysis-method and Appendix LABEL:appendix:glba-extraction for more details). We iteratively refined our prompt until achieving satisfactory performance. We further manually verified and corrected all results. For HTML-based GLBA notices, we parsed them with BeautifulSoup and used regex to identify and extract the three target items, followed by manual verification and correction. Once extracted, we analyzed a GLBA notice’s data-sharing table and opt-out box with regex and manually annotated the OII box.
@@ 4.2.3. Non-GLBA Policy Analysis
General, online, mobile, and CCPA privacy policies rarely disclose the specific data types being shared with third parties and refer to them broadly as “personal information” or “your information”, and occasionally specify a particular data type (e.g., medical information). In contrast, the third-party cookies disclosed in cookie policies are usually associated with data collected during a user’s online activity. Privacy policies use different language to describe third-party sharing practices for marketing. The CCPA defines “sharing” as the disclosure of personal information to a third party for cross-context behavioral advertising, and some CCPA policies adopt this terminology. Other policies, including some CCPA, use broader phrasing (e.g., “we work with advertising companies”) (see more in Appendix LABEL:appendix-clear-loose-examples). In both cases, the practice involves consumer information being disclosed to an entity other than the one that collected it, for use in marketing, though the language used has different regulatory implications.
While existing automated classification methods can identify if a privacy policy segment relates to third-party sharing/collection (e.g., Srinath et al., 2021; Harkous et al., 2018), they are trained on the OPP-115 corpus developed in 2016 (Wilson et al., 2016), which pre-dates CCPA and thus lacks respective notices. Other automated methods (Andow et al., 2019b, 2020; Hooda et al., 2024) extract third-party sharing statements based on pre-defined data type and entity taxonomies, which may lead to incomplete results. Instead, we used a bottom-up manual annotation approach to capture third-party sharing statements as written and respective nuances.
@@ 4.2.4. Manual Annotation Approach
The first author developed an initial codebook drawing from both inductive codes from a preliminary analysis of 60 notices and relevant codes from OPP-115 (Wilson et al., 2016). The codebook was refined over multiple rounds of annotation and discussions among three co-authors on subsets of additional notices (98 in total) across various privacy policy types, structures, formats, and lengths. In each round, the three co-authors independently annotated a set of 5–10 policies, discussed disagreements, and clarified definitions. This process continued until high inter-rater reliability was achieved (Krippendorff’s Cu-$\alpha$ (González-Prieto et al., 2023; Krippendorff, 2019): 0.95 for CCPA privacy policies; 0.89 for general, mobile, and cookie policies).
The final annotation scheme consisted of several code groups that are often applied in combination: affirmation or denial of third-party sharing/selling of personal information, data types involved, sharing purposes, opt-out choice type, and opt-out choice scope. We did not include the receiving entity, as this information was typically vague or not provided. Since privacy policies are often vague or ambiguous (Reidenberg et al., 2016), we included an “unclear” option in each group. We developed distinct (but partially overlapping) codebooks for CCPA-specific and non-CCPA content due to their differing language (see codebooks in Appendix LABEL:appendix:ccpa-codebook and LABEL:appendix:general-codebook), which improved annotation accuracy and IRR. Annotators applied different code combinations to semantically different segments. Given the large volume of privacy policies and the extensive training our annotators underwent to achieve high inter-rater reliability, each policy was annotated once by a single annotator.
@@ 4.2.5. Resulting Dataset
We successfully collected privacy policies from 2,069 (4 lacked policies), and analyzed third-party cookies on 2,070 (3 could not be rendered by our web-browser). We classified the content of all identified privacy policies into five types defined in Section 4.1.1. Among the 2,004 banks for which we found GLBA notices, 2 of them did not include a data-sharing table or disclose sharing practices. 14 banks provided more than one GLBA notice tailored to their different business lines, such as savings, wealth management, and home loans. [^3] For 9 banks, the data-sharing table was consistent across their own GLBA notices. Disclosures on nonaffiliate sharing varied for 5 banks. Opt-out boxes differed only in minor details, such as the exact contact methods, but otherwise were the same for all. For consistency, we analyzed one GLBA notice per bank (notice covering checking/savings services that all banks in our dataset provide). This resulted in a final sample of 2,002 GLBA notices, which our RQ2 GLBA-related results are based on.
@@ 4.3.1. Third-Party Sharing Statements
To identify inconsistencies, we matched up third-party sharing statements in GLBA notices with those in general, mobile, and cookie policies as well as CCPA statements. Notably, these different policy types use different language to describe related third-party sharing practices. To compare disclosures across policies, we identified 15 types of disclosed sharing practices based on our analysis (4 GLBA-specific; 8 in general, mobile, and cookie policies; 3 CCPA-specific). We then analyzed (mis)matches among these sharing practices across a bank’s policy documents. We discuss the sharing practice types, how they match up, and our respective (in)consistency findings in Section 5.2.
@@ 4.4. Limitations
Our study may have several limitations. First, our data collection may not capture all relevant privacy policies due to variation in how banks name and structure these documents on their websites. To maximize coverage, we (1) used an inclusive list of keywords, (2) crawled sites up to three depths deep, and (3) manually inspected sites that our crawler could not access, where key notices (i.e., GLBA) or any privacy-related documents were not detected. Second, annotators’ interpretation of the privacy policies we work with may have led to minor inaccuracies in qualitative coding. To ensure reliability, we developed detailed guidelines, curated a reference list of agreed segments, conducted extensive training, and achieved high inter-rater reliability among three annotators.
@@ 5.1. RQ1: Amount of Information
Table 1. Privacy policies provided by banks (n=2069). Row sums indicate how many banks provide each combination; column sums show totals for each policy type.
@@
GLBA only 1089 1089
GLBA+1 other 406 653
@@
Table 1 summarizes the types of privacy policies provided by each bank and the number of banks offering each combination. We found GLBA notices for 2,004 banks, and 45.3% of all banks provided multiple privacy policies. 915 banks (44.2%) provided one or more additional privacy policies besides GLBA, the most common addition (406, 19.6%) being a general privacy policy, followed by a mobile privacy policy (167, 8.1%). Among banks that provided two or more additional privacy policies besides GLBA notices (262, 12.7%), the most common combination (107, 5.2%) included a GLBA notice, a general privacy policy, and a CCPA-specific notice. This variety demonstrates that for many banks, their privacy disclosures are layered and potentially fragmented across multiple privacy policies.
@@
We used word count and readability as proxies for the amount and complexity of information that a bank’s privacy policies present to consumers. When combining all privacy policies from each bank, total word counts varied widely across banks, with most falling between 554 and 2,192 words ($median$=769, $mean$=1,572). The much higher mean compared to the median reflects a right-skewed distribution driven by a subset of exceptionally lengthy policies (e.g., privacy policy embedded in a large PDF file that also includes statements like online banking agreements). Their readability as measured by FKGL falls mostly within the high school to early college range ($median$=13.2, $mean$=13.6). Given that FKGL scores correspond to U.S. school grade levels, this indicates that most banks’ content requires reading skills well above the average U.S. adult level of 8th grade (Sparx, 2024).
@@
We used bank rank by consolidated assets as a proxy for bank size, where a lower rank indicates a larger bank. We found a modest but statistically significant negative correlation between bank rank and word count ($\rho$=-0.38,$p$$\ll$0.001) and between bank rank and readability score ($\rho$=-0.28,$p$$\ll$0.001), as shown in Figure 3. The largest-sized (i.e. the lowest-ranked banks with consolidated $\geq$\$15,995 million) provided significantly more privacy policy content ($median$=4,017) in contrast to smaller banks offering content generally below 1,500 word count. Larger banks also tend to use more complex language, with FKGL readability scores above 15, and smaller banks average closer to 13 or below.
RQ1 Summary. Both the types of privacy policies and the amount of information vary widely across banks. About half of all banks (936, 45.2%) provided at least two privacy policies, and most banks’ privacy policies are difficult to read. Larger banks provide more privacy policy content with more difficult language compared to smaller banks. While this higher privacy policy content may reflect regulatory pressures or more diverse service offerings, it also raises questions about whether consumers can successfully navigate and comprehend this amount of privacy disclosures.

### v2

@@ 4.1. Data Collection
We used a list of the largest commercial banks by the Federal Reserve (of Governors of the Federal Reserve System, 2024), with consolidated assets exceeding \$300 million, each uniquely identified by its RSSD ID, a Federal Reserve-assigned identifier (nn=2,129). [^1] Consumer-oriented institutions also include savings banks and state non-member banks. They operate at smaller, regional scales and were not included in our study. We use the consolidated asset total as a proxy for customer base and consumer reach: these banks account for 97.3% of all assets held by FDIC-insured commercial banks (52.8% of them by number). We used the FDIC’s BankFind Suite (Federal Deposit Insurance Corporation, 2025) to obtain bank website URLs and branch locations based on their RSSD IDs. We removed 19 duplicate URLs associated with banks owned by the same holding company, [^2] A bank holding company may include multiple legally distinct banks (each with a unique RSSD ID) but typically provides a shared privacy policy across subsidiaries. as well as 19 duplicates where different URLs redirected to the same website due to mergers or acquisitions. We retained only the highest-ranked occurrence in each duplicate case, resulting in 2,091 unique bank website URLs. Because our study includes analysis of banks’ CCPA-related disclosures and opt-out implementations, we collected all data from a California vantage point using a commercial VPN. We successfully accessed 2,073 bank websites, which comprise the final dataset for this study. For each bank, we collected its (1) privacy policies, (2) cookie opt-out controls, (3) CCPA opt-out links, (4) responses to the Global Privacy Control (GPC) signal, and (5) examined third-party cookies. Data collection ran from October 2024 to January 2025. Figure  2 shows our data collection pipeline.
@@ 4.1.1. Privacy policy collection
Privacy policy retrieval.
We built a custom crawler based on Scrapy (Developers, 2008) that searches for links containing a list of privacy policy-related keywords in either the link’s anchor text or URL. Our keyword list was informed by prior research (Amos et al., 2021; Libert, 2018; Wagner, 2023b; Srinath et al., 2021) and iteratively refined and tested on a sample of 100 banks. Our crawler searched for relevant links up to a depth of three to ensure reliable discovery of privacy policies. We manually visited 373 bank websites where the crawler failed to access pages, found no privacy-related content, or no GLBA notices (which we expected all banks to provide). We also manually reviewed the top 200 banks due to more complex site structures. In total, we collected 11,265 potential privacy policy files. Through automated and manual review, we filtered out non-relevant files and duplicates, resulting in 3,305 unique privacy files across 2,069 banks.
@@
Often banks had a webpage titled “Privacy Policy” that contained a collection of privacy policy information, such as a GLBA notice, a CCPA privacy policy and/or a CCPA notice at collection, data practices related to their online and mobile services, and/or a cookie policy. Some banks presented some or all notices in separate, stand-alone policies. To answer RQ1 (number of privacy policies a bank provides), we manually labeled files based on headings and categorized them into five types: (1) GLBA notice, typically titled “U.S. consumer privacy notice,” or “privacy policy.” It is typically presented as a PDF or sometimes as HTML, either standalone or combined with other notices and resembles the GLBA short-form template. (2) General privacy policy, often titled “privacy policy” or “digital privacy” that focuses on online privacy, or may combine elements that would otherwise appear in the following three types of standalone notices: (3) CCPA privacy policy, by which we refer to both the privacy policy and the notice at collection provided in compliance with the CCPA; (4) mobile privacy policy, specific to mobile applications or mobile data collection practices; (5) cookie/advertising policy, focusing on cookie use and tracking technologies or interest-based advertising.
@@ 4.1.3. Third-party cookies
We used third-party cookies as one indicator of whether a bank shares consumer information with external entities. We rendered each site using a Selenium-based web browser. We then used the Chrome DevTools Protocol’s Network domain to monitor network activity and collect all cookies stored during the browser session. Each cookie’s domain was compared to the domain of the website visited; if the domains did not match, the cookie was classified as a third-party cookie. After collecting third-party cookies from each website, we attempted to infer their purpose by comparing their domains against EasyList’s known list of advertising-related trackers (Fanboy et al., 2025). If no exact match was found, we compared only the last two segments of the cookie’s domain with the list (e.g., ”unrulymedia.com” instead of ”.targeting.unrulymedia.com”), which improved detecting advertising-related cookies. We found 1,454 banks (70.1%) allowing third-party cookies on their websites, 1,252 (60.4%) of which contained marketing cookies.
@@ 4.2. Privacy Policy Analysis
To address RQ1 on the amount of privacy information provided by each bank, we extracted the plain text for all of a bank’s policy files. For PDF files, we used a vision-based LLM pipeline (GPT-4o) to extract and compile text from each page, as text extraction methods (e.g, using PyMuPDF (Inc., 2025)) proved unreliable due to the visual complexity of the layouts (e.g., flipped cell texts in GLBA notice tables). For HTML files, we used Boilerpipe (Riebold, 2023) to extract the main text and BeautifulSoup (Richardson, 2025) as a fallback. Then, we merged all of a bank’s processed privacy policy files into one plain-text file. For readability analysis, we used the widely-adopted Flesh-Kincaid Grade Level (Shedlosky-Shoemaker et al., 2009) that measures the difficulty of reading a text based on sentence length and word complexity with a score corresponding to a U.S. school grade.
@@ 4.2.2. GLBA notice analysis
GLBA notices mostly follow the standardized GLBA template, with the majority in PDF or HTML format. We analyzed the data-sharing table, the opt-out box, and the OII box in them (see Figure 1). For PDFs, we employed our vision-based LLM pipeline and refined it given the known GLBA table format: we first manually annotated the page numbers of each target item and then extracted the relevant items from each page into a prescribed format using task-specific prompts. We iteratively refined our prompt until achieving satisfactory performance. We further manually verified and corrected all results. For HTML-based GLBA notices, we parsed them with BeautifulSoup and used regex to identify and extract the three target items, followed by manual verification and correction. Once extracted, we analyzed a GLBA notice’s data-sharing table and opt-out box with regex and manually annotated the OII box.
@@ 4.2.3. Non-GLBA policy analysis
General, online, mobile, and CCPA privacy policies rarely disclose the specific data types being shared with third parties and refer to them broadly as “personal information” or “your information”, and occasionally specify a particular data type (e.g., medical information). In contrast, the third-party cookies disclosed in cookie policies are usually associated with data collected during a user’s online activity. Privacy policies use different language to describe third-party sharing practices for marketing. The CCPA defines “sharing” as the disclosure of personal information to a third party for cross-context behavioral advertising, and some CCPA policies adopt this terminology. Other policies, including some CCPA policies, use broader phrasing (e.g., “we work with advertising companies”). In both cases, the practice involves consumer information being disclosed to an entity other than the one that collected it, for use in marketing, though the language used has different regulatory implications.
While existing automated classification methods can identify if a privacy policy segment relates to third-party sharing/collection (e.g., Srinath et al., 2021; Harkous et al., 2018), they are trained on the OPP-115 corpus developed in 2016 (Wilson et al., 2016), which pre-dates CCPA and thus lacks respective notices. Other automated methods (Andow et al., 2019b, 2020; Hooda et al., 2024) extract third-party sharing statements based on pre-defined data types and entity taxonomies, which may lead to incomplete results. Instead, we used a bottom-up manual annotation approach to capture third-party sharing statements as written and respective nuances.
@@ 4.2.4. Manual annotation approach
The first author developed an initial codebook drawing from both inductive codes from a preliminary analysis of 60 notices and relevant codes from OPP-115 (Wilson et al., 2016). The codebook was refined over multiple rounds of annotation and discussions among three co-authors on subsets of additional notices (98 in total) across various privacy policy types, structures, formats, and lengths. In each round, the three co-authors independently annotated a set of 5–10 policies, discussed disagreements, and clarified definitions in twice-weekly training sessions over eight weeks. After each round, the first author refined the codebook by compiling lists of common and vague phrases and identifying cross-document matches. This process continued until high inter-rater reliability was achieved (Krippendorff’s Cu-$\alpha$ (González-Prieto et al., 2023; Krippendorff, 2019): 0.95 for CCPA privacy policies; 0.89 for general, mobile, and cookie policies).
The final annotation scheme consisted of several code groups that are often applied in combination: affirmation or denial of third-party sharing/selling of personal information, data types involved, sharing purposes, opt-out choice type, and opt-out choice scope. We did not include the receiving entity, as this information was typically vague or not provided. Since privacy policies are often vague or ambiguous (Reidenberg et al., 2016), we included an “unclear” option in each group. We developed distinct (but partially overlapping) codebooks for CCPA-specific and non-CCPA content due to their differing language, which improved annotation accuracy and IRR. Annotators applied different code combinations to semantically different segments. Given the large volume of privacy policies and the extensive training our annotators underwent to achieve high inter-rater reliability, each policy was annotated once by a single annotator. [^3] We provide the collection of documents, codebook, and annotations in a Zenodo repository: https://doi.org/10.5281/zenodo.17014519.
@@ 4.2.5. Resulting dataset
We successfully collected privacy policies from 2,069 bank websites (4 lacked policies), and analyzed third-party cookies on 2,070 websites (3 could not be rendered by our web browser). We classified the content of all identified privacy policies into five types defined in Section 4.1.1. Among the 2,004 banks for which we found GLBA notices, 2 of them did not include a data-sharing table or disclose sharing practices. 14 banks provided more than one GLBA notice tailored to their different business lines, such as savings, wealth management, and home loans. [^4] For 9 banks, the data-sharing table was consistent across their own GLBA notices. Disclosures on nonaffiliate sharing varied for 5 banks. Opt-out boxes differed only in minor details, such as the exact contact methods, but otherwise were the same for all. For consistency, we analyzed one GLBA notice per bank (notice covering checking/savings services that all banks in our dataset provide). This resulted in a final sample of 2,002 GLBA notices, on which our RQ2 GLBA-related results are based.
@@ 4.3.1. Third-Party sharing statements
To identify inconsistencies, we matched up third-party sharing statements in GLBA notices with those in general, mobile, and cookie policies, as well as CCPA statements. Notably, these different policy types use different language to describe related third-party sharing practices. To compare disclosures across policies, we identified 15 types of disclosed sharing practices based on our analysis (4 GLBA-specific; 8 in general, mobile, and cookie policies; 3 CCPA-specific). We then analyzed (mis)matches among these sharing practices across a bank’s policy documents. We discuss the sharing practice types, how they match up, and our respective (in)consistency findings in Section 5.2.
@@ 4.4. Limitations
Our study may have several limitations. First, our data collection may not capture all relevant privacy policies due to variation in how banks name and structure these documents on their websites. To maximize coverage, we (1) used an inclusive list of keywords, (2) crawled sites up to three levels deep, and (3) manually inspected sites that our crawler could not access or where key notices (i.e., GLBA) or any privacy-related documents were not detected. Second, annotators’ interpretation of the privacy policies we analyzed may have led to minor inaccuracies in qualitative coding. To ensure reliability, we developed detailed guidelines, curated a reference list of agreed segments, conducted extensive training, and achieved high inter-rater reliability among three annotators.
@@ 5.1. RQ1: Amount of Information
Table 1. Privacy policies provided by banks (n=2,069). Row sums indicate how many banks provide each combination; column sums show totals for each policy type.
@@
GLBA only 1,089 1,089
GLBA+1 other 407 653
@@
Table 1 summarizes the types of privacy policies provided by each bank and the number of banks offering each combination. We found GLBA notices for 2,004 banks, and 45.2% of all banks provided multiple privacy policies. 915 banks (44.2%) provided one or more additional privacy policies besides their GLBA notice; the most common addition being a general privacy policy (407, 19.6%), followed by a mobile privacy policy (167, 8.1%). Among banks that provided two or more additional privacy policies besides GLBA notices (262, 12.7%), the most common combination (109, 5.3%) included a GLBA notice, a general privacy policy, and a CCPA-specific notice. This variety demonstrates that for many banks, their privacy disclosures are layered and potentially fragmented across multiple privacy policies.
@@
We used word count and readability as proxies for the amount and complexity of information that a bank’s privacy policies present to consumers. When combining all privacy policies from each bank, total word counts varied widely across banks, with most falling between 554 and 2,192 words (medianmedian=769, meanmean=1,572). The much higher mean compared to the median reflects a right-skewed distribution driven by a subset of exceptionally lengthy policies (e.g., privacy policy embedded in a large PDF file that also includes statements like online banking agreements). Their readability as measured by FKGL falls mostly within the high school to early college range (medianmedian=13.2, meanmean=13.6). Given that FKGL scores correspond to U.S. school grade levels, this indicates that most banks’ content requires reading skills well above the average U.S. adult reading level of 8th grade (Sparx, 2024).
@@
We used bank rank by consolidated assets as a proxy for bank size, where a lower rank indicates a larger bank. We found a modest but statistically significant negative correlation between bank rank and word count ($\rho$=-0.38, $p\ll$0.001) and between bank rank and readability score ($\rho$=-0.28, $p\ll$0.001), as shown in Figure 3. The largest-sized (i.e., the lowest-ranked banks with consolidated assets $\geq$\$15,995 million) provided significantly more privacy policy content (medianmedian=4,017) in contrast to smaller banks offering content generally below a 1,500 word count. Larger banks also tend to use more complex language, with FKGL readability scores above 15, and smaller banks average closer to 13 or below.
RQ1 summary. Both the types of privacy policies and the amount of information vary widely across banks. About half of all banks (936, 45.2%) provided at least two privacy policies, and most banks’ privacy policies are difficult to read. Larger banks provide more privacy policy content with more difficult language compared to smaller banks. These results show that despite the GLBA short notice’s intent to simplify privacy disclosures, banks now surround it with multiple, lengthy policies (general, CCPA, cookie), subverting the GLBA notice’s purpose and creating readability/comprehension issues that related work has documented in other sectors.